OpenStudy (anonymous):

Two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively. In how many hours will they meet? A. 1 hr B. 2 hr C. 3 hr

OpenStudy (anonymous):

Together they cover 7 miles each hour. 14 / 7 = 2 hours Algebraically: Let T be the number of hours. 4T + 3T = 14 7T = 14 T = 14/7 T = 2 hours

OpenStudy (anonymous):

Click on best response if you think my answer is good. :)